OpenStudy (anonymous):

I hope someone can help me with this equation. I have to graph f(x)= 2(x-3)^2 + 5 for -1 < or equal to x < or equal to 4.PLEASE HELP!!!

OpenStudy (anonymous):

how far have you gone?

OpenStudy (anonymous):

done*

OpenStudy (anonymous):

i haven't done anything because i am not sure how to draw the graph with "-1 < or equal to x < or equal to 4" on the end

OpenStudy (anonymous):

you know how to draw the graph of y=x^2 right?

OpenStudy (anonymous):

yes, i could draw it if it didn't have the x thing on the end

OpenStudy (anonymous):

you mean the boundaries \[-1\le x \le 4\] ?

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

These just mean that you have to draw the graph only from -1 to 4, so don't continue the curve after these points

OpenStudy (anonymous):